OpenStudy (anonymous):

solve: (x/4)=(x-2/3) A)4 B)8 C)2

OpenStudy (lgbasallote):

\[\frac{x}{4} = x - \frac{2}{3}?\]

OpenStudy (lgbasallote):

\[\frac{x}{4} = x - \frac{2}{3}\] \[\frac{x}{4} - x = -\frac{2}{3}\] \[\frac{x}{4} - \frac{4x}{4} = -\frac{2}{3}\] \[\frac{x-4x}{4} = -\frac{2}{3}\] \[x - 4x = -\frac{2 \times 4}{3}\] \[-3x = -\frac{2 \times 4}{3}\] can you solve for x now?

OpenStudy (anonymous):



OpenStudy (lgbasallote):

\[\frac{x}{4} = \frac{x-2}{3}\] cross multiply \[3x = 4(x-2)\] distribute \[3x = 4x - 8\] can you solve for x now?

OpenStudy (lgbasallote):

do you get it @nique504 ?

OpenStudy (anonymous):

yes. now all i do is add 8 to each side,then divide that answer by 4 right?

OpenStudy (lgbasallote):

no...first put all the x on one side

OpenStudy (lgbasallote):

how can you bring 4x and 3x into one side?

OpenStudy (anonymous):

i can add 4x to both sides?

OpenStudy (lgbasallote):

then that will give you 7x = 8x - 8

OpenStudy (lgbasallote):

right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so what happens now?

OpenStudy (anonymous):

i'm not sure.

OpenStudy (lgbasallote):

remember you want to put the x on one side only so what do you do?

OpenStudy (anonymous):

subtract 4x from each side?

OpenStudy (lgbasallote):

if you subtract 4x from both sides you get 3x - 4x = 4x - 4x - 8 3x - 4x = -8 very good the x are now on one side now what is 3x-4x?

OpenStudy (anonymous):

then my answer would be 8

OpenStudy (lgbasallote):

yup x = 8 would be right

OpenStudy (anonymous):

okay, thanks for helping:)

OpenStudy (lgbasallote):

<tips hat>